# Coefficients-Preserving Sampling for Reinforcement Learning with Flow Matching

### v1

@@ 4.4 Discussion
Inspired by the derivation of VP SDE ( Song et al. (2021b)), which uses Taylor expansion for formula derivation, we can also get an approximate SDE from Flow-CPWS. Note that by Taylor expansion, $\sqrt{t^{2}-x}=t-\frac{x}{2t}+O(x)$ around $x=0$, the above equation can be transformed to,
$\bm{x}_{t-\Delta t}$ $\approx\left(1-(t-\Delta t)\right)\hat{\bm{x}}_{0}+\left(t-\Delta t-\frac{\sigma_{t}^{2}\Delta t}{2(t-\Delta t)}\right)\hat{\bm{x}}_{1}+\sigma_{t}\sqrt{\Delta t}\bm{\epsilon}$
@@
which is the same with Flow-SDE (Equation 9). The approximate equality holds when $\Delta t\to 0$ [^1] The $\sigma_{t}$ controls the noise magnitude, which cannot be very small.. Now we can conclude:
@@
Flow-SDE is a first-order Taylor approximation of Flow-CPWS in the limit of $\Delta t\to 0$, with a noise level error of $\sqrt{\frac{(\sigma_{t}\Delta t)^{2}}{t}+(\frac{\sigma_{t}^{2}\Delta t}{2t})^{2}}$.
@@
For traditional diffusion methods, such as DDPM (Ho et al. (2020)), the sampling step is set as $1000$, so the condition $\Delta t\to 0$ is well satisfied. However, for modern diffusion and flow matching samplers (Song et al. (2021a); Lu et al. (2022)), the sampling step is usually less than $20$. With some distillation techniques (Song et al. (2023); Yin et al. (2024)), the sampling step can be reduced to $4$ or even $1$. The condition $\Delta t\to 0$ no longer holds in these settings. This is the fundamental reason why Flow-SDE produces inaccurate noise levels.
Furthermore, because of the $\frac{1}{t}$ term after Taylor expansion, the approximation error will be huge around $t=0$ (Figure 2), so it is inaccurate even when the sampling step number is high. We provide an alternative to bypass this issue in Appendix B.
@@ 5 Experiments
In this section, we will evaluate the performance of Flow-CPS in the circumstance of GRPO-based reward optimization on four reward models, GenEval (Ghosh et al. (2023)), Text Rendering (OCR Cui et al. (2025)), PickScore (Kirstain et al. (2023)) and HPSv2 (Wu et al. (2023)).
@@ 5.1 Experimental Setup
To make the experiments more convincing, we do experiments on two baselines, Flow-GRPO (Liu et al. (2025)) and Dance-GRPO (Xue et al. (2025)). We follow their experimental settings and only change the sampling method and the log-probability. All the experiments are conducted on $8\timesNVIDIA$ GPUs. We introduce two kinds of tasks, verifiable rewards (RLVR) and preference rewards (RLHF), to evaluate our proposed method.
5.2 Clean Image Sampling
Table 1: GenEval Results on base model Esser et al. (2024) and base code Liu et al. (2025)
@@ 6 Conclusion
This paper utilized LLMs to assist with formula derivations and writing refinement.
@@ Appendix A VP-SDE is an Approximation of DDPM
Similar to Theorem 1, VP-SDE (Song et al. (2021b)) can also be seen as a first-order Taylor approximation of DDPM. In the VP-SDE, the forward process is,
@@
where the approximate equality holds when $\Delta t\ll 1$. Similar to formula 16, it uses Taylor expansion and omits the second and higher order terms.
@@
For the VP SDE backward process, please refer to Appendix E of Song et al. (2021b), which also uses Taylor expansion and omits high-order terms in the derivation. Song et al. (2021b) claimed that the ancestral sampler of DDPM is essentially a discretization of the reverse-time SDE. Conversely, we can also say that the reverse process of the VP-SDE is a continuous approximation of DDPM ancestral sampling.
@@
Figure 7: The ideal and SDE noise level for Equation 18.
Figure 8: Image sampled by Equation 18. There is no obvious noise on these images.
Appendix B An Alternative for the numerical problem
@@
In section 4.4, we mentioned that Flow-SDE has a numerical problem because of the $\frac{1}{t}$ term. One possible patch would be setting $\sigma_{t}=\eta(t-\Delta t)$, then the Flow-SDE becomes,
$\bm{x}_{t-\Delta t}$ $\approx\left(1-(t-\Delta t)\right)\hat{\bm{x}}_{0}+\left(t-\Delta t-\eta^{2}(t-\Delta t)\Delta t\right)\hat{\bm{x}}_{1}+\eta(t-\Delta t)\sqrt{\Delta t}\bm{\epsilon}.$ (18)
@@
Even though this formula still cannot satisfy the requirements of CPS, it has a smaller error than the original Flow-SDE. We show the noise level in Figure 7 and sampled images in Figure 8. It would be useful when the characteristics of the SDE are to be exploited.

### v4

@@ 4.4 Discussion
Inspired by the derivation of VP-SDE ( song2021score), which uses Taylor expansion for formula derivation, we can also get an approximate SDE from Flow-CPWS. Note that by Taylor expansion, $\sqrt{t^{2}-x}=t-\frac{x}{2t}+O(x^{2})$ around $x=0$, the above equation can be transformed to,
$\bm{x}_{t-\Delta t}$ $=\left(1-(t-\Delta t)\right)\hat{\bm{x}}_{0}+\left(t-\Delta t-\frac{\sigma_{t}^{2}\Delta t}{2(t-\Delta t)}+O\left((\sigma_{t}^{2}\Delta t)^{2}\right)\right)\hat{\bm{x}}_{1}+\sigma_{t}\sqrt{\Delta t}\bm{\epsilon}$ (18)
$\approx\left(1-(t-\Delta t)\right)\hat{\bm{x}}_{0}+\left(t-\Delta t-\frac{\sigma_{t}^{2}\Delta t}{2(t-\Delta t)}\right)\hat{\bm{x}}_{1}+\sigma_{t}\sqrt{\Delta t}\bm{\epsilon}$ (19)
@@
[…]
@@
Flow-SDE is a first-order Taylor approximation of Flow-CPWS in the limit of $\sigma_{t}\sqrt{\Delta t}\ll t-\Delta t$ and $\Delta t\to 0$, with a noise level error of $\sqrt{\frac{(\sigma_{t}\Delta t)^{2}}{t}+(\frac{\sigma_{t}^{2}\Delta t}{2t})^{2}}$.
@@
For traditional diffusion methods, such as DDPM (ho2020denoising), the sampling step is set as $1000$, so the condition $\Delta t\to 0$ is well satisfied. However, for modern diffusion and flow matching samplers (songd2021enoising; lu2022dpm), the sampling step is usually less than $20$. With some distillation techniques (song2023consistency; yin2024one), the sampling step can be reduced to $4$ or even $1$. The condition $\Delta t\to 0$ no longer holds in these settings. This is the fundamental reason why Flow-SDE produces inaccurate noise levels.
Furthermore, because of the $\frac{1}{t}$ term introduced by the Taylor expansion, the approximation error will be significant around $t=0$ (Figure 2). Consequently, the approximation remains inaccurate even with a high number of sampling steps. We provide an alternative to mitigate this issue in Appendix C.
@@ 5 Experiments
In this section, we will evaluate the performance of Flow-CPS in the circumstance of GRPO-based reward optimization on four reward models, GenEval (ghosh2023geneval), Text Rendering (OCR cui2025paddleocr), PickScore (kirstain2023pick) and HPSv2 (wu2023human).
@@ 5.1 Experimental Setup
To make the experiments more convincing, we do experiments on two baselines, Flow-GRPO (liu2025flow) and Dance-GRPO (xue2025dancegrpo). We follow their experimental settings and only change the sampling method and the log-probability. All the experiments are conducted on $8\times$ NVIDA A100 GPUs. We introduce two kinds of tasks, verifiable rewards (RLVR) and preference rewards (RLHF), to evaluate our proposed method.
Table 1: GenEval Results on base model esser2024scaling and base code liu2025flow
@@ 5.3 Experimental Results
The baselines reported in Table 1 4 4 4 employ the log-prob definition (Equation 14) from their respective original papers. In contrast, for Flow-CPS, we adopt the formulation in Equation 15. For an ablation study concerning the log-prob definition, please refer to Appendix D.
@@ 6 Conclusion
Reproducibility: Our solution, defined by Equation 13 and 15, can be implemented within 10 lines of code, which are provided in the supplementary material.
@@ Appendix A VP-SDE is an Approximation of DDPM
Similar to Theorem 1, VP-SDE (song2021score) can also be seen as a first-order Taylor approximation of DDPM. In the VP-SDE, the forward process is,
@@
where the approximate equality holds when $\Delta t\ll 1$. Similar to formula 20, it uses Taylor expansion and omits the second and higher order terms. It is interesting to note that Equation 21 is very similar to Equation 20, which also uses Taylor expansion in the first approximation and omits a $\Delta t$ in the second approximation.
@@
For the VP-SDE backward process, please refer to Appendix E of song2021score, which also uses Taylor expansion and omits high-order terms in the derivation. song2021score claimed that the ancestral sampler of DDPM is essentially a discretization of the reverse-time SDE. Conversely, if a pre-trained DDPM is given, we can also say that the reverse process of the VP-SDE is a continuous approximation of DDPM ancestral sampling.
Appendix B DPM-Solver series
DPM-Solver (lu2022dpm) and its variants DPM-Solver++ series(lu2025dpm), also provide SDE solvers. In this section, we will verify if they meet the requirements of CPS.
For SDE-DPM-Solver-1, the sampling function is,
$\bm{x}_{t}=\frac{\alpha_{t}}{\alpha_{s}}\bm{x}_{s}-2\sigma_{t}(e^{h}-1)\hat{\bm{x}}_{1}+\sigma_{t}\sqrt{e^{2h}-1}\bm{\epsilon},$ (22)
where $e^{h}=\frac{\alpha_{t}}{\sigma_{t}}\frac{\sigma_{s}}{\alpha_{s}}$, $\alpha_{t}=1-t$ and $\sigma_{t}=t$ in the concept of Flow Matching. Reformulate it into the factorized form,
$\bm{x}_{t}=\alpha_{t}\hat{\bm{x}}_{0}+\left(\frac{\alpha_{t}}{\alpha_{s}}\sigma_{s}-2\sigma_{t}(e^{h}-1)\right)\hat{\bm{x}}_{1}+\sigma_{t}\sqrt{e^{2h}-1}\bm{\epsilon}.$ (23)
The coefficient of sample is $\alpha_{t}$, which exactly matches the first condition of CPS. However, the total noise level is,
$\sigma_{total}$ $=\sqrt{\left(\frac{\alpha_{t}}{\alpha_{s}}\sigma_{s}-2\sigma_{t}(e^{h}-1)\right)^{2}+\sigma_{t}^{2}(e^{2h}-1)}$
$=\sqrt{\left(\sigma_{t}e^{h}-2\sigma_{t}(e^{h}-1)\right)^{2}+\sigma_{t}^{2}(e^{2h}-1)}$
$=\sigma_{t}\sqrt{(2-e^{h})^{2}+(e^{2h}-1)}$
$=\sigma_{t}\sqrt{2e^{2h}-4e^{h}+3}$
$=\sigma_{t}\sqrt{2(e^{h}-1)^{2}+1}$
$\geq\sigma_{t},$ (24)
where the equality holds only when $e^{h}=1$, so the SDE-DPM-Solver-1 is not Coefficient-Preserving Sampling.
For SDE-DPM-Solver++1, the sampling function is,
$\bm{x}_{t}=\frac{\sigma_{t}}{\sigma_{s}}e^{-h}\bm{x}_{s}+\alpha_{t}(1-e^{-2h})\hat{\bm{x}}_{0}+\sigma_{t}\sqrt{1-e^{-2h}}\bm{\epsilon}.$ (25)
Reformulate it into the factorized form,
$\bm{x}_{t}$ $=\left(\frac{\sigma_{t}}{\sigma_{s}}\alpha_{s}e^{-h}+\alpha_{t}(1-e^{-2h})\right)\hat{\bm{x}}_{0}+\sigma_{t}e^{-h}\hat{\bm{x}}_{1}+\sigma_{t}\sqrt{1-e^{-2h}}\bm{\epsilon}$
$=\left(\alpha_{t}e^{-h}e^{-h}+\alpha_{t}(1-e^{-2h})\right)\hat{\bm{x}}_{0}+\sigma_{t}e^{-h}\hat{\bm{x}}_{1}+\sigma_{t}\sqrt{1-e^{-2h}}\bm{\epsilon}$
$=\alpha_{t}\hat{\bm{x}}_{0}+\sigma_{t}e^{-h}\hat{\bm{x}}_{1}+\sigma_{t}\sqrt{1-e^{-2h}}\bm{\epsilon}.$ (26)
The coefficient of sample is $\alpha_{t}$, which exactly matches the first condition of CPS. The total noise level is,
$\sigma_{total}=\sigma_{t}\sqrt{e^{-2h}+1-e^{-2h}}=\sigma_{t}.$ (27)
Thus, the SDE-DPM-Solver++1 perfectly matches the requirements of CPS. As also verified in lu2025dpm, it is a special case of DDIM with $\eta=\sigma_{t}\sqrt{1-e^{-2h}}$. Our proposed Flow-CPS can be seen as a special case of DDIM with $\eta=\sigma_{t}\sin(\frac{\eta\pi}{2})$ , which retains the hyper-parameter $\eta$ to tune the injected noise level.  [^2] Here we swap the $\eta$ and $\sigma_{t}$ in the main text to follow the mathematical notations in DPM-Solver++.
For higher-order DPM-Solvers, the high-order terms are residuals of two successive estimations of noise or sample, such as $\sigma_{t}(e^{h}-1)\frac{\epsilon_{\theta}(\bm{x}_{r},r)-\epsilon_{\theta}(\bm{x}_{s},s)}{r_{1}}$. Since the coefficients of the two estimations cancel each other out, our analysis above remains unaffected for the higher-order DPM-Solvers.
Figure 7 illustrates the training curves for both Flow-CPS and SDE-DPM-Solver++1 utilizing the OCR reward. We observe that Flow-CPS becomes unstable when $\eta\leq 0.5$, characterized by intermittent and sudden drops in reward. Similarly, the training curves for SDE-DPM-Solver++1 exhibit comparable sudden drops, which typically correlate with a lack of diversity. In the right panel of Figure 7, we plot the injected noise level of SDE-DPM-Solver++1 alongside the equivalent $\eta$ in Flow-CPS. Notably, the equivalent $\eta$ for SDE-DPM-Solver++1 varies across timesteps: it initiates at $1.0$, gradually decreases to approximately $0.43$, and subsequently returns to $1.0$. The underlying cause of the significant instability observed in SDE-DPM-Solver++1 remains under investigation and requires further research.
This training instability leads to inconsistency in the final performance. As shown in Table 5, we conducte two separate experiments using SDE-DPM-Solver++1. The peak rewards for these runs vary from $0.966$ to $0.970$, highlighting the variance in outcomes. Such instability is detrimental to reproducibility and necessitates multiple trials to obtain a satisfactory model.
Table 5: OCR Results for SDE-DPM-Solver++1 and Flow-CPS
Model DPM++ run1 DPM++ run2 CPS $\eta=0.3$ CPS $\eta=0.5$ CPS $\eta=0.7$
Reward 0.966 0.970 0.973 0.963 0.975
Figure 7: Left: The training curves of Flow-CPS with the OCR reward. Middle: The training curves of SDE-DPM-Solver++1 with the OCR reward. Right: The equivalent $\eta$ value for SDE-DPM-Solver++1.
Appendix C An Alternative for the numerical problem
@@
Figure 8: The ideal and SDE noise level for Equation 29. The error is ignorable when $\eta\leq 1$ for 44 and more steps.
@@
In section 4.4, we mentioned that Flow-SDE has a numerical problem because of the $\frac{1}{t}$ term. Considering the limit of $\sigma_{t}\sqrt{\Delta t}\ll t-\Delta t$ and $\Delta t\to 0$ in Theorem 1, one possible patch would be setting $\sigma_{t}=\eta(t-\Delta t)$. Based on the Formula 18, our modified reverse Flow-SDE becomes,
$\bm{x}_{t-\Delta t}$ $\approx\left(1-(t-\Delta t)\right)\hat{\bm{x}}_{0}+\left(t-\Delta t-\frac{\eta^{2}}{2}(t-\Delta t)\Delta t\right)\hat{\bm{x}}_{1}+\eta(t-\Delta t)\sqrt{\Delta t}\bm{\epsilon}$ (28)
$=\bm{x}_{t}-\hat{\bm{v}}_{\theta}(\bm{x}_{t},t)\Delta t-\frac{\eta^{2}}{2}(t-\Delta t)\Delta t\hat{\bm{x}}_{1}+\eta(t-\Delta t)\sqrt{\Delta t}\bm{\epsilon}$
$\approx\bm{x}_{t}-\hat{\bm{v}}_{\theta}(\bm{x}_{t},t)\Delta t-\frac{\eta^{2}t}{2}\hat{\bm{x}}_{1}\Delta t+\eta t\sqrt{\Delta t}\bm{\epsilon}.$ (29)
The noise level of Equation 28 is $(t-\Delta t)\sqrt{1+(\frac{\eta^{2}\Delta t}{2})^{2}}$, which is slightly higher than $(t-\Delta t)$. In the limit of $\Delta t\to 0$, Equation 29 converges to the following reverse SDE,
$\mathrm{d}\bm{x}_{t}=\hat{\bm{v}}_{\theta}(\bm{x}_{t},t)\mathrm{d}t+\frac{\eta^{2}t}{2}\hat{\bm{x}}_{1}(\bm{x}_{t},t)\mathrm{d}t+\eta t\mathrm{dw}.$ (30)
@@
Even though this formula still cannot meet the requirements of CPS, it has a smaller error than the original Flow-SDE. We show the noise level in Figure 8 and sampled images in Figure 9. It would be useful when the characteristics of the SDE are necessary.
Figure 9: Image sampled by Equation 28 with $\eta=1$. There is no obvious noise on these images.
Appendix D Ablation on the logprob
In Equation 15, we removed the denominator $2\sigma_{t}^{2}$ to prevent numerical instability caused by division by near-zero values in the final diffusion steps. We also applied this modification to the Flow-GRPO baseline for an ablation study. As shown in Figure 10, although this change initially accelerates convergence, the final performance is comparable to the original version.
Figure 10: The ablation on the log-probability. Our algorithm fails to converge with the denominator $2\sigma_{t}^{2}$, so it is not shown in this figure.
Appendix E Qualitative Results
Figures 11 and 12 show visualizations of images optimized by the PickScore and HPSv2 reward models, respectively. Honestly speaking, a higher reward score does not necessarily equate to superior image quality. Often, the optimized images contain an excessive amount of detail, a phenomenon that can be seen as a way to ”hack” the reward model. In practice, a balance must be found between achieving a high reward score and maintaining the image’s visual coherence.
Appendix F The Use of Large Language Models (LLMs)
We utilize LLMs to assist with formula derivations and writing refinement on this paper.
Figure 11: Images created by FLUX.1-dev baseline, Flow-GRPO and Flow-CPS (ours) using PickScore as the reward model. The figures suggest that the PickScore reward model tends to add texture details on the images.
Figure 12: Images created by FLUX.1-dev baseline, Dance-GRPO and Flow-CPS (ours) using HPSv2 as the reward model. The figures suggest that the HPSv2 reward model appears to improve the high-frequency details and the rendering of light and shadow.